# Covariate balancing estimation and model selection for difference-in-differences approach

### v1

@@ 3.1 Estimation method and statistical properties
From these equations, ${\rm E}[d^{[1]}\bm{x}\bm{x}^{{\rm T}}]={\rm E}[\{e^{[1]}(\bm{x};\bm{\alpha})/e^{[0]}(\bm{x};\bm{\alpha})\}d^{[0]}\bm{x}\bm{x}^{{\rm T}}]$ is obtained, which means that the expected value of $\bm{x}\bm{x}^{{\rm T}}$ in the treatment group is adjusted to match the expected value of $\bm{x}\bm{x}^{{\rm T}}$ in the control group through weighting by the propensity score. In the original covariate balancing (Imai and Ratkovic 2014), the first-order moments of the covariates are balanced, that is, the $\bm{x}\bm{x}^{{\rm T}}$ in the expectation is naturally set to $\bm{x}$. In this article, the second-order moment of the covariates is balanced.
The semi-vectorization of the upper triangular part of the matrix is written as ${\rm vech}(\cdot)$. Since $\bm{x}\bm{x}^{{\rm T}}$ is a symmetric matrix and defining
@@
the moment condition can be written as
${\rm E}\{\bm{h}(\bm{d},\bm{x};\bm{\alpha})\}=\bm{0}.$ (9)
@@
$\frac{1}{n}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}})=\bm{0},$ (10)
@@
$\frac{1}{n}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}})\approx\bm{0}$ (11)
is used. Specifically, $\hat{\bm{\alpha}}$ can be obtained by the generalized method of moments (Hansen 1982) or the generalized empirical likelihood method (Qin and Lawless 1994).
Here, we use the generalized method of moments and propose $\hat{\bm{\alpha}}^{\rm CB}\equiv\mathop{\rm argmin}_{\bm{\alpha}}\{\bm{h}_{n}(\bm{\alpha})^{{\rm T}}\bm{W}_{n}\bm{h}_{n}(\bm{\alpha})\}$ as the estimator of $\bm{\alpha}$ by the Covariate balancing (CB) method, where $\bm{h}_{n}(\bm{\alpha})\equiv 1/n\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})$. Defining $\bm{G}_{n}(\bm{\alpha})\equiv\partial\bm{h}_{n}(\bm{\alpha})/\partial\bm{\alpha}$, $\hat{\bm{\alpha}}^{\rm CB}$ satisfies
Here, $\bm{W}_{n}$ is an arbitrary semi-positive definite matrix, the simplest of which is the identity matrix, but the optimal matrix, $\{1/n\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})^{{\rm T}}\}^{-1}$, provides the most efficient estimation of $\bm{\alpha}$. However, this optimal matrix is not always practical because it may be too close to the degenerate matrix depending on the data, causing instability in estimation.
@@
The idea of using constraints that approximately satisfy the empirical version of the moment condition to estimate the parameters of the propensity score has also been used in Ning et al. (2020b). In that article, dealing with the case of high dimensional covariates, they propose a doubly robust estimation method that combines the semiparametric estimation method of ATT with the sparse estimation method. It is also assumed that the outcome is expressed in a linear model of covariates, and to improve the convergence speed of the Horvitz-Thompson estimator, the weak covariate balancing property:
$\frac{1}{n}\sum_{i=1}^{n}\biggl{(}\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha})}-1\biggr{)}\bm{\alpha}^{{\rm T}}\bm{x}_{i}\approx 0$
is included as a constraint in estimating the parameter of the propensity score.
Here, we state the regularity conditions of the generalized method of moments (Newey and McFadden 1994), which are necessary to describe the properties of the proposed estimation method.
Assumption 5 (Consistency).
$(\mathrm{i})$ The $\bm{W}_{n}$ converges to $\bm{W}$, which is a semi-positive definite matrix, in probability. Also, $\bm{\alpha}=\bm{\alpha}^{\dagger}$ if and only if ${\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})]=\bm{0}$. $(\mathrm{ii})$ The $\bm{\alpha}$ is the interior of the compact set $\bm{A}\subset\mathbb{R}^{p}$. $(\mathrm{iii})$ The $\bm{h}(\bm{d},\bm{x};\bm{\alpha})$ is differentiable with probability 1 for each $\bm{\alpha}\in\bm{A}$, that is, $e^{[1]}(\bm{x};\bm{\alpha})$ is differentiable with probability 1. $(\mathrm{iv})$ ${\rm E}[\sup_{\bm{\alpha}\in\bm{A}}\|\bm{h}(\bm{d},\bm{x};\bm{\alpha})\|]<\infty$.
Assumption 6 (Asymptotic normality).
$(\mathrm{i})$ $\bm{\alpha}^{\dagger}$ is interior of $\bm{A}$. $(\mathrm{ii})$ $\bm{h}(\bm{d},\bm{x};\bm{\alpha})$ is differentiable in a neighborhood $\Gamma$ of $\bm{\alpha}^{\dagger}$, with probability approaching one. $(\mathrm{iii})$ ${\rm E}[\|\bm{h}(\bm{d},\bm{x};\bm{\alpha}^{\dagger})\|^{2}]<\infty$. $(\mathrm{iv})$ ${\rm E}[\sup_{\bm{\alpha}\in\Gamma}\|\partial\bm{h}(\bm{d},\bm{x};\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}\|_{\rm F}]<\infty$. $(\mathrm{v})$ $\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\allowbreak\bm{G}(\bm{\alpha}^{\dagger})$ is nonsingular for $\bm{G}(\bm{\alpha})\equiv{\rm E}[\partial\bm{h}(\bm{d},\bm{x};\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}]$.
Under the above regularity conditions, the proposed method has the following double robustness.
Theorem 1.
Replacing the propensity score in (7) by $e^{[1]}(\bm{x};\hat{\bm{\alpha}}^{\rm CB})$, we name the covariate balancing for DID (CBD) estimator and denote it by $\hat{\bm{\theta}}^{\rm CBD}$. If the propensity score model $e^{[1]}(\bm{x};\bm{\alpha})$ is correctly specified, or if the change in the outcome follows a linear model of covariates, that is, if
${\rm E}[\Delta^{[k]}\mid\bm{x},d^{[k]}=1]=\bm{x}^{{\rm T}}\bm{\beta}^{[k]*}$ (13)
holds, then under Assumptions 1-6, this $\hat{\bm{\theta}}^{\rm CBD}$ will converge to $\bm{\theta}^{*}$ in probability. Namely, proposed estimator has double robustness.
Proof.
From Assumption 5, $\hat{\bm{\alpha}}^{\rm CB}$ is consistent with $\bm{\alpha}^{\dagger}$. By expanding $\bm{h}_{n}(\hat{\bm{\alpha}})$ around $\bm{\alpha}^{\dagger}$ and applying Assumption 6 and (12), it is shown to be
$\sqrt{n}(\hat{\bm{\alpha}}^{\rm CB}-\bm{\alpha}^{\dagger})$
=$=-\{\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{G}_{n}(\hat{\bm{\alpha}}^{\rm CB})\}^{-1}\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\sqrt{n}\bm{h}_{n}(\bm{\alpha}^{\dagger})+{\rm o}_{\rm P}(1)$ (14)
where $\bm{\Omega}(\bm{\alpha})\equiv{\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})\bm{h}^{{\rm T}}(\bm{d},\bm{x};\bm{\alpha})]$. From the consistency of $\hat{\bm{\alpha}}^{\rm CB}$,
$\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*}$
=$=\bigg{\{}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg{\}}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}$
=$=\bigg{\{}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg{\}}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}+{\rm o}_{\rm P}(1)$ (15)
is also obtained. If the propensity score model is correct and $\bm{\alpha}^{\dagger}=\bm{\alpha}^{*}$ is satisfied, that is, if $e^{[1]}(\bm{x})=e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})$ and $e^{[0]}(\bm{x})=e^{[0]}(\bm{x};\bm{\alpha}^{\dagger})$ hold, then from (5), $\hat{\bm{\theta}}^{\rm CBD}$ is a consistent estimator for $\bm{\theta}^{*}$. On the other hand, even if the model for the propensity score is not correct, when the change in outcome follows a linear model of covariates, that is, when (7) holds,
is obtained from (3). Then, from (6),
$\bm{\theta}^{*}={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]=\bm{\beta}^{[1]*}-\bm{\beta}^{[0]*}$
holds. Therefore,
$={\rm E}\biggl{[}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bigg{\{}\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}{\rm E}[\Delta_{i}^{[1]}\mid\bm{x}_{i},d^{[1]}_{i}=1]-\frac{d^{[0]}_{i}}{e^{[0]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}{\rm E}[\Delta_{i}^{[0]}\mid\bm{x}_{i},d^{[0]}_{i}=1]-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\bigg{\}}\biggr{]}$
$={\rm E}\biggl{[}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\biggl{\{}\biggl{(}\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}-1\biggr{)}\bm{x}_{i}^{{\rm T}}\bm{\beta}^{*}_{1}-\biggl{(}\frac{d^{[0]}_{i}}{e^{[0]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}-1\biggr{)}\bm{x}_{i}^{{\rm T}}\bm{\beta}^{*}_{0}\biggr{\}}\biggr{]}$
$={\rm E}\biggl{[}\bm{H}_{1}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{\beta}^{*}_{1}-\bm{H}_{0}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{\beta}^{*}_{0}\biggr{]}$
$=\bm{0}$
is obtained. Here, the last equality holds from the moment condition (9). From the above, when the model of the propensity score is correctly specified, or when the change in outcomes follows a linear model of covariates,
holds in (15) and $\hat{\bm{\theta}}^{\rm CBD}$ is consistent. ∎
Remark 2.
Here, we discuss the relationship between the proposed method and recent methods. Sant’Anna and Zhao (2020) proposed a doubly robust estimating method in the study of estimating ATT by SDID, in which the propensity score and outcome are estimated by parametric models, respectively. If one of the models is correctly specified, their proposed estimator is consistent with the ATT. On the other hand, our approach does not estimate the parameters of the model for the outcomes, although the estimator is doubly robust because of the use of covariate balancing. In addition, the second-order moment of the covariates is balanced across groups targeting the treatment group population.

### v3

@@ 3.1 Estimation method and its statistical properties
From these conditions is obtained ${\rm E}[d^{[1]}\bm{x}\bm{x}^{{\rm T}}]={\rm E}[\{e^{[1]}(\bm{x};\bm{\alpha})/e^{[0]}(\bm{x};\bm{\alpha})\}d^{[0]}\bm{x}\bm{x}^{{\rm T}}]$, which means that the expectation of $\bm{x}\bm{x}^{{\rm T}}$ in the treatment group is adjusted to match the expectation of $\bm{x}\bm{x}^{{\rm T}}$ in the control group through weighting based on the propensity scores. In the original covariate balancing, the first-order moments of the covariates are balanced; that is, $\bm{x}\bm{x}^{{\rm T}}$ in the expectation is naturally set to $\bm{x}$. In this paper, the second-order moments of the covariates are balanced.
The semi-vectorization of the upper triangular part of the matrix is written as ${\rm vech}(\cdot)$. Since $\bm{x}\bm{x}^{{\rm T}}$ is a symmetric matrix, by letting
@@
the moment conditions can be rewritten as
${\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})]=\bm{0}.$ (8)
We will denote an $\bm{\alpha}$ satisfying (8) as $\bm{\alpha}^{\dagger}$. Since the propensity score model may be misspecified, $\bm{\alpha}^{\dagger}$ is not necessarily the true value $\bm{\alpha}^{*}$, but when the propensity score model is correctly specified, $\bm{\alpha}^{\dagger}=\bm{\alpha}^{*}$. In practice, the SDID approach uses a $\hat{\bm{\alpha}}$ satisfying the empirical version of the moment conditions,
$\frac{1}{n}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}})=\bm{0}.$ (9)
If the moment conditions are just-identified, that is, the dimension of the moment conditions (the number of constraints) is the same as the dimension of the parameter $\bm{\alpha}$, then $\bm{\alpha}$ that satisfies (9) exists. On the other hand, if they are over-identified, that is, the dimension of the moment conditions is larger than the dimension of $\bm{\alpha}$, then an $\hat{\bm{\alpha}}$ that satisfies (9) does not necessarily exist. In that case, an $\hat{\bm{\alpha}}$ satisfying
$\frac{1}{n}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}})\approx\bm{0}$
is used. Specifically, $\hat{\bm{\alpha}}$ is obtained from the generalized method of moments (Hansen 1982) or the generalized empirical likelihood method (Qin and Lawless 1994).
Here, we will use the generalized method of moments and propose $\hat{\bm{\alpha}}^{\rm CB}\equiv\mathop{\rm argmin}_{\bm{\alpha}}\{\bm{h}_{n}(\bm{\alpha})^{{\rm T}}\allowbreak\bm{W}_{n}\allowbreak\bm{h}_{n}(\bm{\alpha})\}$ as the estimator of $\bm{\alpha}$ by using the covariate balancing (CB) method, where $\bm{h}_{n}(\bm{\alpha})\equiv n^{-1}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})$ and $\bm{W}_{n}$ is an arbitrary semi-positive definite matrix. Upon defining $\bm{G}_{n}(\bm{\alpha})\equiv\partial\bm{h}_{n}(\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}$, one sees that the estimator $\hat{\bm{\alpha}}^{\rm CB}$ satisfies
Its simplest candidate is the identity matrix, but the optimal matrix, $\{n^{-1}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})\allowbreak\bm{h}(\bm{d}_{i},\allowbreak\bm{x}_{i};\bm{\alpha})^{{\rm T}}\}^{-1}$, provides the most efficient estimator of $\bm{\alpha}$. However, this optimal matrix is not necessarily practical since it may be too close to being a degenerate matrix for some data, causing considerable instability in the estimation of $\bm{\alpha}$ and consequently $\bm{\theta}$.
Remark 1.
The idea of using constraints that approximately satisfy the empirical version of the moment conditions to estimate the parameters of the propensity scores has also been examined. In Ning et al. (2020b) a doubly robust estimation method is proposed for dealing with the case of high-dimensional covariates. This method combines a semiparametric estimation of the ATT with sparse estimation. It is assumed that the outcome can be expressed by a linear model of covariates, and to speed up the convergence of the Horvitz-Thompson estimator, the weak covariate balancing property, written as
$\frac{1}{n}\sum_{i=1}^{n}\bigg\{\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha})}-1\bigg\}\bm{\alpha}^{{\rm T}}\bm{x}_{i}\approx 0,$
is included as a constraint in estimating the parameter of the propensity scores.
Here, we state the regularity conditions of the generalized method of moments (Newey and McFadden 1994) that are necessary to derive the properties of the proposed estimation method.
Assumption 5 (Consistency).
$(\mathrm{i})$ $\bm{W}_{n}$ converges in probability to $\bm{W}$, which is a semi-positive definite matrix. Also, $\bm{\alpha}=\bm{\alpha}^{\dagger}$ if and only if $\bm{W}{\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})]=\bm{0}$. $(\mathrm{ii})$ $\bm{\alpha}^{\dagger}$ is in the interior of the compact set $\bm{A}\subset\mathbb{R}^{p}$. $(\mathrm{iii})$ $\bm{h}(\bm{d},\bm{x};\bm{\alpha})$ is continuous; that is, $e^{[1]}(\bm{x};\bm{\alpha})$ is continuous at each $\bm{\alpha}\in\bm{A}$ with probability one. $(\mathrm{iv})$ ${\rm E}[\sup_{\bm{\alpha}\in\bm{A}}\|\bm{h}(\bm{d},\bm{x};\bm{\alpha})\|]<\infty$.
Assumption 6 (Asymptotic normality).
$(\mathrm{i})$ $\bm{h}(\bm{d},\bm{x};\bm{\alpha})$ is continuously differentiable in a neighborhood $\Gamma$ of $\bm{\alpha}^{\dagger}$ with probability approaching one. $(\mathrm{ii})$ ${\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha}^{\dagger})]=\bm{0}$ and ${\rm E}[\|\bm{h}(\bm{d},\bm{x};\bm{\alpha}^{\dagger})\|^{2}]\allowbreak<\infty$. $(\mathrm{iii})$ ${\rm E}[\sup_{\bm{\alpha}\in\Gamma}\allowbreak\|\partial\bm{h}(\bm{d},\bm{x};\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}\|_{\rm F}]<\infty$. $(\mathrm{iv})$ $\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\allowbreak\bm{G}(\bm{\alpha}^{\dagger})$ is non-singular for $\bm{G}(\bm{\alpha})\equiv{\rm E}[\partial\bm{h}(\bm{d},\bm{x};\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}]$.
Under these regularity conditions, the proposed method can be shown to be doubly robust, as follows. Here, we name the estimator obtained by replacing the propensity scores with $e^{[1]}(\bm{x};\hat{\bm{\alpha}}^{\rm CB})$ in (7) the covariate balancing for DID (CBD) estimator and denote it as $\hat{\bm{\theta}}^{\rm CBD}$.
Theorem 1.
If the propensity score model $e^{[1]}(\bm{x};\bm{\alpha})$ is correctly specified, or if the change in outcomes follows a linear model of covariates, that is, if
${\rm E}[\Delta^{[k]}\mid\bm{x},d^{[k]}=1]=\bm{x}^{{\rm T}}\bm{\beta}^{[k]*}$ (11)
holds, then under Assumptions 1-6, $\hat{\bm{\theta}}^{\rm CBD}$ converges in probability to $\bm{\theta}^{*}$. That is, the proposed estimator has double robustness.
Proof.
@@
$\sqrt{n}(\hat{\bm{\alpha}}^{\rm CB}-\bm{\alpha}^{\dagger})$
$=-\{\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{G}_{n}(\hat{\bm{\alpha}}^{\rm CB})\}^{-1}\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\sqrt{n}\bm{h}_{n}(\bm{\alpha}^{\dagger})+{\rm o}_{\rm P}(1)$ (12)
where $\bm{\Omega}(\bm{\alpha})\equiv{\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})\bm{h}(\bm{d},\bm{x};\bm{\alpha})^{{\rm T}}]$. From the consistency of $\hat{\bm{\alpha}}^{\rm CB}$, we also obtain
$\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*}$
$=\bigg\{\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg\}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}$ (13)
$=\bigg\{\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg\}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}+{\rm o}_{\rm P}(1).$ (14)
If the propensity score model is correctly specified and $\bm{\alpha}^{\dagger}=\bm{\alpha}^{*}$ is satisfied, that is, if $e^{[1]}(\bm{x})=e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})$ and $e^{[0]}(\bm{x})=e^{[0]}(\bm{x};\bm{\alpha}^{\dagger})$ hold, then from (5), $\hat{\bm{\theta}}^{\rm CBD}$ is a consistent estimator of $\bm{\theta}^{*}$. On the other hand, even if the propensity score model is misspecified, when the change in outcomes follows a linear model of covariates, that is, when (7) holds, we obtain
from (3). Then, it follows from (6) that
$\bm{\theta}^{*}={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]=\bm{\beta}^{[1]*}-\bm{\beta}^{[0]*}.$
Therefore, we obtain
$={\rm E}\bigg[e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bigg\{\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}{\rm E}[\Delta_{i}^{[1]}\mid\bm{x}_{i},d^{[1]}_{i}=1]$
@@
$={\rm E}\bigg[e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bigg[\bigg\{\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}-1\bigg\}\bm{x}_{i}^{{\rm T}}\bm{\beta}^{*}_{1}-\bigg\{\frac{d^{[0]}_{i}}{e^{[0]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}-1\bigg\}\bm{x}_{i}^{{\rm T}}\bm{\beta}^{*}_{0}\bigg]\bigg]$
$={\rm E}[\bm{H}_{1}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{\beta}^{*}_{1}-\bm{H}_{0}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{\beta}^{*}_{0}]=\bm{0}.$
Here, the last equality holds from the moment conditions in (8). From the above, when the propensity score model is correctly specified, or when the change in outcomes follows a linear model of covariates, it holds that
in (14), and therefore $\hat{\bm{\theta}}^{\rm CBD}$ is consistent. ∎
@@
Let us discuss the relationship between our proposed method and the method in Sant’Anna and Zhao (2020). Unlike our setting, which estimates the conditional ATT, the latter directly estimates the ATT using a doubly robust approach. In their approach as well, the propensity scores are estimated so that the first-order moments of the distribution for the covariates are balanced. They also suppose a parametric model of the covariates for the change in the control group’s outcomes and estimate it, using these estimators to estimate the ATT. In contrast, we place importance on evaluating the heterogeneity of the ATT, including variable selection, and then set a function of the covariates for the ATT. We suppose a parametric model of the covariates for the change in outcomes as well, although not estimating it. Then, if the propensity scores are estimated so that not the first-order but the second-order moments of the distribution for the covariates are balanced, it is easily, yet somewhat unexpectedly, shown that the conditional ATT estimator is doubly robust.